# Modelling bulk mechanical effects in a planar cellular monolayer

### v1

@@ Abstract
We use a three-dimensional formulation of the cell vertex model to describe the mechanical properties of a confluent planar monolayer of prismatic cells. Treating cell height as a degree of freedom, we reduce the model to a two-dimensional form. We show how bulk effects, associated with cell volume and total surface area, lead to coupling between energy variations arising from changes in cell apical area and apical perimeter, a feature missing from standard implementations of the two-dimensional vertex model. The model identifies four independent mechanisms by which cells can lose in-plane rigidity, relating to variations in total cell surface area, the strength of lateral adhesion, and constrictive forces at the apical cortex. The model distinguishes bulk from in-plane stresses, and identifies two primary measures of cell shear stress. In the rigid regime, the model shows how lateral crowding in a disordered isolated monolayer can lead to cell elongation towards the monolayer centre. We examine loss of in-plane rigidity in a disordered monolayer and connect isolated patches of stiffness that persist during the rigidity transition to the spectrum of a Laplacian matrix. This approach enables bulk mechanical effects in an epithelium to be captured within a two-dimensional framework.
@@ I Introduction
Unlike traditional hyperelasticity, in which mechanical energy is described in terms of strains, the vertex model assigns an energy to a multicelluar tissue that depends on geometric invariants (lengths, areas and volumes of individual cells). The vertex model is now routinely implemented in order to investigate the organisation of tissues in 3D [2], facilitated by the release of computational tools [3, 4]. In comparison to its standard 2D formulation, in which energy is typically assumed to be the sum of convex functions of cell apical area and apical perimeter, 3D vertex models offer a broader range of constitutive choices. The energy may include a contribution that is linear in contact area between adherent cells [5, 6], linear in the area exposed to surroundings [5, 7], linear in the area between heterotypic cells [8, 9], quadratic in cell volume relative to a target value $V_{0}^{*}$ [5, 10], quadratic in total cell surface area relative to a target area $a_{0}^{*}$ [10], or quadratic in cell edge lengths [11]; see [12] for an overview of some of the different combinations assumed by prior authors. Additional energy contributions may penalise bending of cell edges [13] or bending of adjacent cells within a sheet [14].
Beyond providing the opportunity to infer stresses in an epithelium from primarily geometric information [15, 16], the 2D vertex model has attracted considerable attention as a tool with which to investigate rigidity transitions in planar monolayers [17]. In the standard 2D framework, rigidity of individual cells arises from a balance between forces that resist changes in cell apical area and forces that resist changes in the apical perimeter. The latter forces regulate resistance to in-plane shear, enabling the rigidity of a monolayer to be regulated through changes of a single dimensionless parameter. A rigidity transition in a 3D multicellular tissue with energy defined in terms of cell volume and total surface area has been shown to be regulated by the parameter $a_{0}^{*}/V_{0}^{*2/3}$ [10, 18, 19]. In the configuration that we investigate below, using a 3D description of a planar monolayer, we identify four independent dimensionless parameters that may regulate the in-plane rigidity transition, broadening the potential repertoire of mechanisms by which cells may exploit this change of phase.
@@
In this study, we consider a planar monolayer of cells of uniform (or near-uniform) thickness, seeking to derive a 2D model that retains 3D mechanical features. To facilitate this, we impose apical-basal symmetry, postponing investigation of more complex structural features such as scutoids [28, 29] and cell nuclei. We pose an energy that is quadratic in cell volume, total area, apical and basal perimeter, and linear in adherent area between neighbouring cells (features illustrated in Fig. 1). We use a force balance to determine the thickness of the cell monolayer (in contrast to others [29] who prescribe the thickness), imposing a constraint that restricts height variations between neighbouring cells. We disregard curvature of the monolayer, recognising its importance in many applications [30, 31], in order to focus on the implications of our chosen constitutive model. Our approach is similar in spirit to [7] and [32], who also reduced 3D models of epithelial sheets to 2D formulations: however we disregard cell-substrate adhesion, which allows for a discontinuous squamous/columnar transition [7]; and we do not adopt a phase-field approach [32]. We compare the reduced 2D model to the traditional 2D vertex model in which apical area and apical perimeter contributions make independent contributions to the total energy; despite having four independent dimensionless parameters at our disposal, we find no direct reduction that leads to the classical model. We examine rigidity transitions as functions of 3D parameters for a monolayer of identical hexagonal prisms. For monolayers that are spatially disordered, we investigate relationships between 2D and 3D stress components, discuss rigidity in terms of connected regions of stiffness, and demonstrate crowding-induced elongation of cells at the centre of an isolated monolayer in the rigid regime. The model is laid out in Sec. II and numerical results are presented in Sec. III.
@@ II.1 Prismatic cells
In what follows, we index cells with $h=1,\dots,N_{c}$, faces with $i=1,\dots,N_{f}$, edges with $j=1,\dots,N_{e}$ and vertices with $k=1,\dots,2N_{v}$. The normals assigned to apical and basal faces of the monolayer are assumed to be parallel to the constant vector $\hat{\mathbf{z}}$. Cell shapes are specified in terms of vertex locations $\mathbf{R}^{\pm}_{k}\in\mathbb{R}^{3}$ and the depth $H$. We assume vertices $\mathbf{R}_{k}^{-}$, for $k=N_{v}+1,\dots,2N_{v}$ lie on the plane $z=0$ and vertices $\mathbf{R}_{k}^{+}$, for $k=1,\dots,N_{v}$, lie on the plane $z=H$. Edges and faces are indexed in the order apical ($+$), lateral ($l$) and basal (--), so that $N_{e}=N_{e}^{+}+N_{e}^{l}+N_{e}^{-}$ and $N_{f}=N_{f}^{+}+N_{f}^{l}+N_{f}^{-}$. Writing $\mathbf{R}_{k}^{-}=\mathbf{r}_{k}^{-}$ for $k=N_{v}+1,\dots,2N_{v}$ and $\mathbf{R}_{k}^{+}=\mathbf{r}_{k}^{+}+H\hat{\mathbf{z}}$ for $k=1,\dots,N_{v}$, where $\hat{\mathbf{z}}\cdot\mathbf{r}_{k}^{\pm}\equiv 0$, we evaluate geometric quantities in terms of vertex locations $\mathbf{r}_{k}^{\pm}$, expressing these quantities in terms of signed incidence and adjacency matrices (see Appendix A).
@@
A dynamic implementation of the cell vertex model involves $12N_{v}+$ 1 ordinary differential equations (ODEs) describing the evolution of $\mathsfbf{r}^{\pm}(t)$ and $H(t)$ down the energy gradients identified in (3) to an equilibrium state. Assuming the apical and basal vertices share the same topological organisation, this state is likely to evolve to one in which $\mathbf{r}_{k}^{+}=\mathbf{r}_{k+N_{v}}^{-}$ for $k=1,\dots,N_{v}$.
We therefore assume apical-basal symmetry, allowing us to write $\mathbf{r}_{k}^{+}=\mathbf{r}_{k+N_{v}}^{-}\equiv\mathbf{r}_{k}$, where $\mathbf{r}_{k}$ is a vector confined to a plane orthogonal to $\hat{\mathbf{z}}$. This simplifies the geometric quantities substantially, allowing them to be expressed in terms of apical areas and perimeters as $V_{h}=V_{h}(A_{h}^{+},H)$, $a_{h}=a_{h}(A_{h}^{+},L_{h}^{+},H)$ and $A^{l}_{i}=A^{l}_{i}(t_{i}^{+},H)$, where $t_{i}^{+}$ identifies the apical edge associated with lateral face $i$. Specifically,
@@ II.2 A constitutive model
Finally, to reduce the number of parameters, we scale $a_{0}$ from the equilibrium problem by setting
$\begin{gathered}A_{h}^{+}=a_{0}\tilde{A}_{h},~{}H=\frac{\tilde{H}}{a_{0}},~{}L_{h}^{+}=a_{0}^{2}\tilde{L}_{h},~{}P=\frac{\tilde{P}}{a_{0}},~{}T=\frac{\tilde{T}}{a_{0}^{2}},\\ \Gamma_{a}=\frac{\tilde{\Gamma}_{a}}{a_{0}^{2}},~{}\Gamma_{A}=\frac{\tilde{\Gamma}_{A}}{a_{0}},~{}\Gamma_{L}=\frac{\tilde{\Gamma}_{L}}{a_{0}^{4}},~{}L_{0}=a_{0}^{2}\tilde{L}_{0}.\end{gathered}$ (13)
This allows us to write (10) and (12) as (dropping tildes)
@@ II.3 Monolayer configurations
For an array of identical hexagonal prismatic cells of edge-length $l$, $A_{h}\equiv A=\tfrac{1}{2}3\sqrt{3}l^{2}$, $L_{h}\equiv L=6l$ and the horizontal force (11a) becomes
$a_{0}\frac{\partial U}{\partial\mathbf{r}_{k}^{+}}={\color[rgb]{0,0,0}2}\sum_{h}\left\{3\sqrt{3}lP^{2D}_{h}+6T^{2D}_{h}\right\}\frac{\partial l}{\partial\mathbf{r}_{k}^{+}}.$ (15)
@@
00 0 $=2AP^{2D}(A,L,H)+LT^{2D}(A,L,H),$ (16a)
$A$ $=\tfrac{1}{2}3\sqrt{3}l^{2},\quad L=6l,$ (16b)
@@
Elimination of $H$, $A$ and $T$ from (14a,b) and (16) yields a single nonlinear scalar equation, with roots specifying possible values of $l$. Some asymptotic limits of configurations for rigid hexagons are described in Appendix E. As in the 2D model (Appendix C; Fig. 10 below), physical solutions are limited to certain regions of parameter space. We restrict attention to solutions with $l>0$ and $H>0$; where multiple equilibria exist, we consider only solutions that are stable when vertices move under a frictional drag, as determined by analysis of eigenvalues of the relevant Hessian (Appendix F). We expect solutions for which $T^{2D}>0$ to be rigid.
@@
$\frac{L}{\sqrt{A}}\geq s_{n}=2\sqrt{n\tan\left(\frac{\pi}{n}\right)},$ (18)
@@
In addition to reporting numerical solutions of (17), and where appropriate of (19), plus relevant asymptotic limits (Appendix E), we also report distributions of the shape parameter $L_{h}/\sqrt{A_{h}}$ for isolated disordered monolayers, to reveal the nature of the rigidity transition. To determine these, we used the VertexModel.jl package [34] to simulate 500-cell monolayers. Monolayers with stress-free external boundary conditions were grown with stochastic cell divisions, allowing T1 transitions to occur when edge lengths fell below a threshold of 0.01 [35], resulting in a disordered system. We replaced the tension and pressure in the 2D vertex model (14a,b) with $T^{2D}$ and $P^{2D}$, enforcing the vertical force balance (14c). Monolayers were relaxed to equilibrium with a tolerance (in vertex location) of $5\times 10^{-6}$. Simulations are reported in Section II.4 below.
@@ II.5 Tissue-level stress and external loading
$V_{G}\bm{\sigma}_{G}\equiv{\textstyle{\sum_{h}}}g_{h}V_{h}{\bm{\sigma}_{h}}={\textstyle{\sum_{h}}}g_{h}\left(A_{h}{\bm{\sigma}_{h}^{2D}}+\frac{\partial U_{h}}{\partial H}H\hat{\mathbf{z}}\otimes\hat{\mathbf{z}}\right).$ (28)
@@
$2A\frac{\partial\mathcal{U}}{\partial A}+L\frac{\partial\mathcal{U}}{\partial L}=2AHP_{\mathrm{ext}},~{}~{}\frac{\partial\mathcal{U}}{\partial H}=0,~{}~{}A=\frac{\sqrt{3}}{24}L^{2}.$ (30)
In contrast, isotropic external compression of a monolayer of near-identical hexagons via ${\bm{\sigma}_{G}}=P_{\mathrm{ext}}\mathsf{I}_{3}$ leads to the constraints $P^{\mathrm{eff},2D}=HP_{\mathrm{ext}}$ and $\partial\mathcal{U}/\partial H=AP_{\mathrm{ext}}$, requiring modification of (16a) and (16c). This is equivalent to the system
$2A\frac{\partial\mathcal{U}}{\partial A}+L\frac{\partial\mathcal{U}}{\partial L}=2AHP_{\mathrm{ext}},~{}~{}\frac{\partial\mathcal{U}}{\partial H}=AP_{\mathrm{ext}},~{}~{}A=\frac{\sqrt{3}}{24}L^{2}.$ (31)
@@ III.1 Regular hexagonal cells: parameter variation
Volume $V_{h}$ 0.49 Bulk pressure $P_{h}$0.51
@@
Apical area $A_{h}$ 0.24 Cortical tension $T^{c}_{h}$0.03
@@
Monolayer height $H$ 2.01 2D pressure $P^{2D}_{h}$0.39
@@
Table 1: Summary of approximate values of geometric and mechanical quantities for hexagonal cells in the base case, with $\Gamma_{a}=\Gamma_{L}=0.1,\Gamma_{A}=-0.5$ and $L_{0}=2$.
@@
We choose a base case where $\Gamma_{a}=\Gamma_{L}=0.1,\Gamma_{A}=-0.5$ and $L_{0}=2$, yielding geometric and mechanical variables listed in Table 1. This example is in the rigid regime, with positive tension, $T^{2D}_{h}$, balancing negative pressure, $P^{2D}_{h}$; negative cortical tension $T_{h}^{c}$ is offset by positive surface tension $T_{h}$. We found only one solution for this set of parameters. We examine variations of each parameter in turn before taking a wider view of parameter space. In the 2D quadratic model [1], hexagons lose rigidity when the target apical perimeter exceeds $s_{6}\approx 3.72$ (Appendix C). In the present 3D model, other factors contribute to $T_{h}^{2D}$, and thus to the rigidity transition.
@@
Figure 2: Impact of variation of $L_{0}$. (a) Cell apical area, $A_{h}$, perimeter, $L_{h}$, and height, $H$ as functions of $L_{0}$. Symbols correspond to the cell shapes illustrated in (b). The base case cell shape is shown in grey (solid symbol); cells for $L_{0}={1,2.5}$ are marked by open symbols. (c) Bulk pressure, $P_{h}$ (cyan), bulk tension, $T_{h}$ (green), cortical line tension, $T_{h}^{c}$ (orange), contribution of adhesive tension from a cell’s lateral face, $T_{hi}^{l}=\frac{1}{2}\Gamma_{A}|C_{hi}^{l}|$ (pink), (d) $P^{2D}_{h}$ (blue), $T^{2D}_{h}$ (red) versus $L_{0}$. In all panels, black points at $L_{0}=2$ indicate the base case, the grey shaded region indicates the floppy regime, the hatched regions indicate no physical solutions, and $\Gamma_{a}=\Gamma_{L}=0.1,\Gamma_{A}=-0.5$.
@@
Figure 2 shows how the cell shape, pressures and tensions change with $L_{0}$, with other parameters held at their base-case values. In the rigid regime, increasing $L_{0}$ leads, intuitively, to an increase in apical area and perimeter and a decrease in height (Fig. 2), making cells more squamous. As shown in Fig. 2(d), $T^{2D}$ and $P^{2D}$ fall to zero at $L_{0}\approx 2.83$, resulting in a loss of in-plane rigidity. Beyond the rigidity transition, the apical area (of distorted floppy cells) decreases, going to zero at $L_{0}\approx 7.0$; the perimeter and height continue to increase and decrease respectively. In contrast, reducing $L_{0}$ constricts the apical face until, at $L_{0}\approx-1.26$, the equilibrium area and perimeter vanish. Thus rigid and floppy solutions exist only over a finite range of $L_{0}$ values, and highly elongated (columnar) cells are readily obtained by a reduction in $L_{0}$.
@@
Figure 3: Impact of variation of $\Gamma_{a}$, presented following the format of Fig. 2. (b) The base case cell shape at $\Gamma_{a}=0.1$ is shown in grey (solid symbol); cells for $\Gamma_{a}={0.05,0.5}$ are marked by open symbols. In all panels, $\Gamma_{L}=0.1,\Gamma_{A}=-0.5,L_{0}=2.0$. The dashed lines in (a,c) show the asymptote (45) as $\Gamma_{a}\rightarrow\infty$.
@@
$\Gamma_{a}$ controls the strength of the bulk tension associated with variations of total surface area. This term contributes to the compound pressure and tension terms at the apical surface, and is responsible for coupling $A_{h}$, $L_{h}$ and $H$ in determining $P^{2D}$ and $T^{2D}$. Figure 3 shows the impact of changing $\Gamma_{a}$. In the rigid regime, increasing $\Gamma_{a}$ causes the cell to become more cuboidal, consistent with expectation of increasing the overall surface tension. In-plane rigidity can be lost either by decreasing or, less intuitively, increasing $\Gamma_{a}$ past thresholds. As $\Gamma_{a}$ increases, the more squamous floppy solution asymptotes quickly towards the limiting case of fixed surface area (45), which corresponds here to $L=2=L_{0},H=1/2,A=0$ (achieved by strong in-plane distortions). When $\Gamma_{a}\rightarrow 0$ the cell area in the more columnar floppy state shrinks to zero, $P_{h}$ and $T_{h}^{c}$ go to zero and the bulk and adhesive tensions, $T_{h}^{c}$ and $T_{hi}=\tfrac{1}{2}\Gamma_{A}|C_{hi}^{l}|$, balance each other.

### v2

@@ Abstract
We use a three-dimensional formulation of the cell vertex model to describe the mechanical properties of a confluent planar monolayer of prismatic cells. Treating cell height as a degree of freedom, we reduce the model to a two-dimensional form. We show how bulk effects, associated with cell volume and total surface area, lead to coupling between energy variations arising from changes in cell apical area and apical perimeter, a feature missing from standard implementations of the two-dimensional vertex model. The model identifies five independent mechanisms by which cells can lose in-plane rigidity, relating to variations in total cell surface area, the strength of lateral adhesion, and constrictive forces at the apical cortex. The model distinguishes bulk from in-plane stresses, and identifies two primary measures of cell shear stress. In the rigid regime, the model shows how lateral crowding in a disordered isolated monolayer can lead to cell elongation towards the monolayer centre. We examine loss of in-plane rigidity in a disordered monolayer and connect isolated patches of stiffness that persist during the rigidity transition to the spectrum of a Laplacian matrix. This approach enables bulk mechanical effects in an epithelium to be captured within a two-dimensional framework.
@@ I Introduction
Unlike traditional hyperelasticity, in which mechanical energy is described in terms of strains, the vertex model assigns an energy to a multicelluar tissue that depends on geometric invariants (lengths, areas and volumes of individual cells). The vertex model is now routinely implemented in order to investigate the organisation of tissues in 3D [2], facilitated by the release of computational tools [3, 4]. In comparison to its standard 2D formulation, in which energy is typically assumed to be the sum of convex functions of cell apical area and apical perimeter, 3D vertex models offer a broader range of constitutive choices. The energy may include a contribution that is linear in contact area between adherent cells [5, 6, 7], linear in the area exposed to surroundings [5, 6, 8], linear in the area between heterotypic cells [9, 10], quadratic in cell volume relative to a target value $V_{0}^{*}$ [5, 6, 11], quadratic in total cell surface area relative to a target area $a_{0}^{*}$ [11], or quadratic in cell edge lengths [12]; see [13] for an overview of some of the different combinations assumed by prior authors. Additional energy contributions may penalise bending of cell edges [14] or bending of adjacent cells within a sheet [15].
Beyond providing the opportunity to infer stresses in an epithelium from primarily geometric information [16, 17], the 2D vertex model has attracted considerable attention as a tool with which to investigate rigidity transitions in planar monolayers [18]. In the standard 2D framework, rigidity of individual cells arises typically from a balance between forces that resist changes in cell apical area and forces that resist changes in the apical perimeter. The latter forces regulate resistance to in-plane shear, enabling the rigidity of a monolayer to be regulated through changes of a single dimensionless parameter. A rigidity transition in a 3D multicellular tissue with energy defined in terms of cell volume and total surface area has been shown to be regulated by the parameter $a_{0}^{*}/V_{0}^{*2/3}$ [11, 19, 20]. In the configuration that we investigate below, using a 3D description of a planar monolayer, we identify five independent dimensionless parameters that may regulate the in-plane rigidity transition, broadening the potential repertoire of mechanisms by which cells may exploit this change of phase.
@@
In this study, we consider a planar monolayer of cells of uniform (or near-uniform) thickness, seeking to derive a 2D model that retains 3D mechanical features. To facilitate this, we impose apical-basal symmetry, postponing investigation of more complex structural features such as scutoids [29, 30] and cell nuclei. We pose an energy that is quadratic in cell volume, total area, apical and basal perimeter, and linear in adherent area between neighbouring cells (features illustrated in Fig. 1). Imposing a constraint that restricts height variations between neighbouring cells, we use a force balance to determine the thickness of the cell monolayer (in contrast to others [30, 31] who prescribe the thickness). We disregard curvature of the monolayer, recognising its importance in many applications [32, 33], in order to focus on the implications of our chosen constitutive model. Our approach is similar in spirit to [8], [34] and [31], who also reduced 3D models of epithelial sheets to 2D formulations: we disregard cell-substrate adhesion [8] and we do not adopt a phase-field approach [34]; but like [8] and [31], we recover a sharp squamous/columnar transition in some circumstances. We compare the reduced 2D model to the traditional 2D vertex model in which apical area and apical perimeter contributions make independent contributions to the total energy; despite having five independent dimensionless parameters at our disposal, we find no direct reduction that leads to the classical model unless cell height is prescribed. We examine rigidity transitions as functions of 3D parameters for a monolayer of identical hexagonal prisms. For monolayers that are spatially disordered, we investigate relationships between 2D and 3D stress components, discuss rigidity in terms of connected regions of stiffness, and demonstrate crowding-induced elongation of cells at the centre of an isolated monolayer in the rigid regime. The model is laid out in Sec. II and numerical results are presented in Sec. III.
@@ II.1 Prismatic cells
In what follows, we index cells with $h=1,\dots,N_{c}$, faces with $i=1,\dots,N_{f}$, edges with $j=1,\dots,N_{e}$ and vertices with $k=1,\dots,2N_{v}$. The normals assigned to apical and basal faces of the monolayer are assumed to be parallel to the constant vector $\hat{\mathbf{z}}$. Cell shapes are specified in terms of vertex locations $\mathbf{R}^{\pm}_{k}\in\mathbb{R}^{3}$ and the depth $H$. We assume vertices $\mathbf{R}_{k}^{-}$, for $k=N_{v}+1,\dots,2N_{v}$ lie on the plane $z=0$ and vertices $\mathbf{R}_{k}^{+}$, for $k=1,\dots,N_{v}$, lie on the plane $z=H$. Edges and faces are indexed in the order apical (++), lateral ($l$) and basal (-), so that $N_{e}=N_{e}^{+}+N_{e}^{l}+N_{e}^{-}$ and $N_{f}=N_{f}^{+}+N_{f}^{l}+N_{f}^{-}$. Writing $\mathbf{R}_{k}^{-}=\mathbf{r}_{k}^{-}$ for $k=N_{v}+1,\dots,2N_{v}$ and $\mathbf{R}_{k}^{+}=\mathbf{r}_{k}^{+}+H\hat{\mathbf{z}}$ for $k=1,\dots,N_{v}$, where $\hat{\mathbf{z}}\cdot\mathbf{r}_{k}^{\pm}\equiv 0$, we evaluate geometric quantities in terms of vertex locations $\mathbf{r}_{k}^{\pm}$, expressing these quantities in terms of signed incidence and adjacency matrices (see Appendix A).
@@
A dynamic implementation of the cell vertex model involves $2N_{v}+1$ ordinary differential equations (ODEs) describing the evolution of $\mathsfbf{r}^{\pm}(t)$ and $H(t)$ down the energy gradients identified in (3) to an equilibrium state.
We then assume apical-basal symmetry, allowing us to write $\mathbf{r}_{k}^{+}=\mathbf{r}_{k+N_{v}}^{-}\equiv\mathbf{r}_{k}$, where $\mathbf{r}_{k}$ is a vector confined to a plane orthogonal to $\hat{\mathbf{z}}$. This simplifies the geometric quantities substantially, allowing them to be expressed in terms of apical areas and perimeters as $V_{h}=V_{h}(A_{h}^{+},H)$, $a_{h}=a_{h}(A_{h}^{+},L_{h}^{+},H)$ and $A^{l}_{i}=A^{l}_{i}(t_{i}^{+},H)$, where $t_{i}^{+}$ identifies the apical edge associated with lateral face $i$. Specifically,
@@ II.2 A constitutive model
Finally, we scale $a_{0}$ from the equilibrium mechanical problem by setting
$\begin{gathered}A_{h}^{+}=a_{0}\tilde{A}_{h},~H=\frac{\tilde{H}}{a_{0}},~L_{h}^{+}=a_{0}^{2}\tilde{L}_{h},~{\color[rgb]{0,0,0}P_{h}^{2D}}=\frac{{\color[rgb]{0,0,0}\tilde{P}_{h}^{2D}}}{a_{0}},\\ {\color[rgb]{0,0,0}T_{h}^{2D}}=\frac{{\color[rgb]{0,0,0}\tilde{T}_{h}^{2D}}}{a_{0}^{2}},\,\Gamma_{a}=\frac{\tilde{\Gamma}_{a}}{a_{0}^{2}},\,\Gamma_{A}=\frac{\tilde{\Gamma}_{A}}{a_{0}},\,\Gamma_{L}=\frac{\tilde{\Gamma}_{L}}{a_{0}^{4}},\,L_{0}=a_{0}^{2}\tilde{L}_{0}.\end{gathered}$ (13)
$a_{0}$ is retained in the geometric ratio $L_{h}^{+}/\sqrt{A_{h}^{+}}=a_{0}^{3/2}\tilde{L}_{h}/\sqrt{\tilde{A}_{h}}$, which takes the value $2^{3/2}3^{1/4}$ for perfect hexagons. Dropping tildes, (13) allows us to write (10) and (12) as
@@ II.3 Monolayer configurations
For an array of identical hexagonal prismatic cells of edge-length $l$, $A_{h}\equiv A=\tfrac{1}{2}3\sqrt{3}{\color[rgb]{0,0,0}a_{0}^{3}}l^{2}$, $L_{h}\equiv L=6l$ and the horizontal force (11a) becomes
$a_{0}\frac{\partial U}{\partial\mathbf{r}_{k}^{+}}={\color[rgb]{0,0,0}2}\sum_{h}\left\{3\sqrt{3}{\color[rgb]{0,0,0}a_{0}^{3}}lP^{2D}_{h}+6T^{2D}_{h}\right\}\frac{\partial l}{\partial\mathbf{r}_{k}^{+}}.$ (15)
@@
$0$ $=2AP^{2D}(A,L,H)+LT^{2D}(A,L,H),$ (16a)
$A$ $=\tfrac{1}{2}3\sqrt{3}{\color[rgb]{0,0,0}a_{0}^{3}}l^{2},\quad L=6l,$ (16b)
@@
Elimination of $H$, AA and $T$ from (14a,b) and (16) yields a single nonlinear scalar equation, with roots specifying possible values of $l$. Some asymptotic limits of configurations for rigid hexagons are described in Appendix E. As in the 2D model (Appendix C; Fig. 10 below), physical solutions are limited to certain regions of parameter space. We restrict attention to solutions with $l>0$ and $H>0$; where multiple equilibria exist, we consider only solutions that are stable when vertices move under a frictional drag, as determined by analysis of eigenvalues of the relevant Hessian (Appendix F). We expect solutions for which $T^{2D}>0$ to be rigid.
@@
$\frac{{\color[rgb]{0,0,0}a_{0}^{3/2}}L}{\sqrt{A}}\geq s_{n}=2\sqrt{n\tan\left(\frac{\pi}{n}\right)},$ (18)
@@
In addition to reporting numerical solutions of (17), and where appropriate of (19), plus relevant asymptotic limits (Appendix E), we also report distributions of the shape parameter $L_{h}/\sqrt{A_{h}}$ for isolated disordered monolayers, to reveal the nature of the rigidity transition. To determine these, we used the VertexModel.jl package [36] to simulate 500-cell monolayers. Isolated monolayers with stress-free external boundary conditions were grown with stochastic cell divisions, taking place uniformly randomly across the monolayer, allowing T1 transitions to occur when edge lengths fell below a threshold of 0.01 [37], resulting in a disordered system. Once $N_{c}=500$, we suppressed divisions and the monolayer was relaxed to the nearest equilibrium. In these simulations, we replaced the tension and pressure in the 2D vertex model (14a,b) with $T^{2D}$ and $P^{2D}$, enforcing the vertical force balance (14c). Monolayers were relaxed to equilibrium with a tolerance (in vertex location) of $5\times 10^{-6}$. Simulations are reported in Section II.4 below.
@@ II.5 Tissue-level stress and external loading
[…]
@@
$2A\frac{\partial\mathcal{U}}{\partial A}+L\frac{\partial\mathcal{U}}{\partial L}=2AHP_{\mathrm{ext}},~~\frac{\partial\mathcal{U}}{\partial H}=0,~~A=\frac{\sqrt{3}}{24}{\color[rgb]{0,0,0}a_{0}^{3}}L^{2}.$ (30)
In contrast, isotropic external compression of a monolayer of near-identical hexagons via ${\bm{\sigma}_{G}}=P_{\mathrm{ext}}\mathsf{I}_{3}$ leads to the constraints $P^{\mathrm{eff},2D}=HP_{\mathrm{ext}}$ and $\partial\mathcal{U}/\partial H=AP_{\mathrm{ext}}$, requiring modification of (16a) and (16c). This is equivalent to the system (30) but with $\partial U/\partial H=0$ replaced by
$\frac{\partial\mathcal{U}}{\partial H}=AP_{\mathrm{ext}}.$ (31)
@@ III.1 Regular hexagonal cells: parameter variation
Volume $V_{h}$ 0.49 Bulk pressure $P_{h}$ -0.51
@@
Apical area $A_{h}$ 0.24 Cortical tension $T^{c}_{h}$ -0.03
@@
Monolayer height $H$ 2.01 2D pressure $P^{2D}_{h}$ -0.39
@@
Table 1: Summary of approximate values of geometric and mechanical quantities for hexagonal cells in the baseline case, with $\Gamma_{a}=\Gamma_{L}=0.1,\Gamma_{A}=-0.5$, $L_{0}=2$ and $a_{0}=1$.
@@
We choose a baseline case where $\Gamma_{a}=\Gamma_{L}=0.1,\Gamma_{A}=-0.5$, $L_{0}=2$ and $a_{0}=1$, yielding geometric and mechanical variables listed in Table 1. This example is in the rigid regime, with positive tension, $T^{2D}_{h}$, balancing negative pressure, $P^{2D}_{h}$; negative cortical tension $T_{h}^{c}$ is offset by positive surface tension $T_{h}$. We found only one solution for this set of parameters. We examine variations of each parameter in turn before taking a wider view of parameter space. In the 2D quadratic model [1], hexagons lose rigidity when the target apical perimeter exceeds $s_{6}\approx 3.72$ (Appendix C). In the present 3D model, other factors contribute to $T_{h}^{2D}$, and thus to the rigidity transition.
@@
Figure 2: Impact of variation of $L_{0}$. (a) Cell apical area, $A_{h}$, perimeter, $L_{h}$, and height, $H$ as functions of $L_{0}$. Symbols correspond to the cell shapes illustrated in (b). The baseline cell shape is shown in grey (solid symbol); cells for $L_{0}={1,2.5}$ are marked by open symbols. (c) Bulk pressure, $P_{h}$ (cyan), bulk tension, $T_{h}$ (green), cortical line tension, $T_{h}^{c}$ (orange), contribution of adhesive tension from a cell’s lateral face, $T_{hi}^{l}=\frac{1}{2}\Gamma_{A}|C_{hi}^{l}|$ (pink), and (d) $P^{2D}_{h}$ (blue), $T^{2D}_{h}$ (brown) versus $L_{0}$. In all panels, black points at $L_{0}=2$ indicate the baseline case, the grey shaded region indicates the floppy regime, the hatched regions indicate no physical solutions, and $\Gamma_{a}=\Gamma_{L}=0.1,\Gamma_{A}=-0.5$, $a_{0}=1$.
@@
Figure 2 shows how the cell shape, pressures and tensions change with $L_{0}$, with other parameters held at their baseline values. In the rigid regime, increasing $L_{0}$ leads, intuitively, to an increase in apical area and perimeter and a decrease in height (Fig. 2a), making cells more squamous. As shown in Fig. 2(d), $T^{2D}$ and $P^{2D}$ fall to zero at $L_{0}\approx 2.83$, resulting in a loss of in-plane rigidity. Beyond the rigidity transition, the apical area (of distorted floppy cells) decreases, going to zero at $L_{0}\approx 7.0$; the perimeter and height continue to increase and decrease respectively. In contrast, reducing $L_{0}$ constricts the apical face until, at $L_{0}\approx-1.26$, the equilibrium area and perimeter vanish. Thus rigid and floppy solutions exist only over a finite range of $L_{0}$ values, and highly elongated (columnar) cells are readily obtained by a reduction in $L_{0}$.
@@
Figure 3: Impact of variation of $\Gamma_{a}$, presented following the format of Fig. 2. (b) The baseline cell shape at $\Gamma_{a}=0.1$ is shown in grey (solid symbol); cells for $\Gamma_{a}={0.05,0.5}$ are marked by open symbols. In all panels, $\Gamma_{L}=0.1,\Gamma_{A}=-0.5,L_{0}=2.0$ and $a_{0}=1$. The dashed lines in (a,c) show the asymptote (47) as $\Gamma_{a}\rightarrow\infty$.
@@
$\Gamma_{a}$ controls the strength of the bulk tension associated with variations of total surface area. This term contributes to the compound pressure and tension terms at the apical surface, and is responsible for coupling $A_{h}$, $L_{h}$ and $H$ in determining $P^{2D}$ and $T^{2D}$. Figure 3 shows the impact of changing $\Gamma_{a}$. In the rigid regime, increasing $\Gamma_{a}$ causes the cell to become more cuboidal, consistent with expectation of increasing the overall surface tension. In-plane rigidity can be lost either by decreasing or, less intuitively, increasing $\Gamma_{a}$ past thresholds. As $\Gamma_{a}$ increases, the more squamous floppy solution asymptotes quickly towards the limiting case of fixed surface area (45, 47), achieved by strong in-plane distortions. When $\Gamma_{a}\rightarrow 0$ the cell area in the more columnar floppy state shrinks to zero, $P_{h}$ and $T_{h}^{c}$ go to zero and the bulk and adhesive tensions, $T_{h}^{c}$ and $T_{hi}=\tfrac{1}{2}\Gamma_{A}|C_{hi}^{l}|$, balance each other.